OpenStudy (anonymous):

Let z1= 2+3i, z2=2-3i, and z3= 4+5i find: z1z2

OpenStudy (anonymous):

does that mean multiply ?

OpenStudy (anonymous):

YAA but complex multiplication but keep 1 thing in mind i^2=-1

OpenStudy (anonymous):

ie i=sqrt(-1)

OpenStudy (anonymous):

i=sqrt(-1) always like that or ?

OpenStudy (anonymous):

always it is iota it appears when there is negative sign inside sqrt likesqrt(-4)=sqrt(4)sqrt(-1)=2i

OpenStudy (anonymous):

so answer is 65

OpenStudy (anonymous):

cuz i^2=-1

OpenStudy (anonymous):

i mean 13

OpenStudy (anonymous):

what r u finding???

OpenStudy (anonymous):

opps 65 was the problem number my bad lol

OpenStudy (anonymous):

Note that this is zz', where z' is the conjugate, so the answer should be a real number. 13 is right.

OpenStudy (anonymous):

yes 13

OpenStudy (anonymous):

yeah my bad lol

OpenStudy (anonymous):

In general, zz' = |z|^2 :-)

OpenStudy (anonymous):

ok thanks bmp ^^